The sum of two numbers is 10. What are the two numbers if the sum of their squares is a minimum?
Answer by mohanraj04(12)

ans:
0+10=10 and (0^2)+(10^2)=100
1+9=10 and (1^2)+(9^2)=82
2+8=10 and (2^2)+(8^2)=68
3+7=10 and (3^2)+(7^2)=58
4+6=10 and (4^2)+(6^2)=52
5+5=10 and (5^2)+(5^2)=50
6+4=10 and (6^2)+(4^2)=52
7+3=10 and (7^2)+(3^2)=58
8+2=10 and (8^2)+(2^2)=68
9+1=10 and (9^2)+(1^2)=82
10+0=10 and (10^2)+(0^2)=100
here the minimum is 50 but it is the sum of same number.
the next minimum is 52 so the two numbers are 4 and 6